dy Solve the initial value problem (t+1). dt = y + (4t² + 4t) (t + 1), y(1) = 9 g(t) =
Answer #1

1 da = 4+ (+++++) (++) y (1)=9 can re-write Difterential Equation So) We ( 4++++) dy + P(H) Y = 8(t) which is (t+1) dy at Solq 4tdttc = ht १ +277 -0 2 (7+123+ (+1+8 = T = b = (1 h use in@ 9 = 2 (1+1) +C (1+1) = 9 = 4+2 e (= = 572 So G (+)= 9+?(1+t) +